At what temperature on the Fahrenheit scale will temperature on a centigrade scale be three times that on a Fahrenheit scale?
1
Expert's answer
2020-06-10T18:22:26-0400

Let, Temperature in Fahrenheit scale is T,thus in centigrade scale temperature would be 3T.

We know that

"\frac{C}{5}=\frac{F-32}{9}"

where , C is temperature in centigrade scale and F is in Fahrenheit scale.


Thus, from above formula we get,

"\frac{3T}{5}=\frac{T-32}{9}\implies 27T=5T-160\\
\implies 22T=-160\implies T=-7.27^{\circ}F"